Identify the variable
If the line that passes through (5,d) and (9,2) 1/2 the slope of -3/4 find the equation of the line associated with the point.
How can I find the line associated with the point especially when I don't have a number and the number is replaced with a letter?

Answer by richwmiller(17219)

What is the formula for finding the slope when you know two points on the line?
fill out the formula with what you know.
m =(y-y1)/(x-x1)
m is the slope
the equation of the line would be
y - y 1 = m( x - x1 )